OpenStudy (anonymous):

Which of the following is NOT equivalent to the rational expression? (Down Below)

OpenStudy (anonymous):

\[\frac{3}{9-x}\] \[A: -\frac{ 3 }{ x-9 }\] \[B: -\frac{ 3 }{ 9-x }\] \[C: \frac{ -3 }{ x-9 }\]

OpenStudy (anonymous):

Choice A is going to be equivalent. It takes the original and multiplies the bottom by negative one. -(9-x)=(x-9). Then they multiplied the top by negative 1 as well by putting the negative in front. Choice B just multiplies the top by negative one, so it's not equivalent. You have to multiply BOTH by negative 1 in order for it to stay the same. What do you think about Choice C?